The volume of a shipping box in a cubic feet can be expressed as the polynomial 2x^3+11x^2+17x+6. Each dimension of the box can be expressed as a linear equations with integer coefficients. Which expression could represent one of the three dimensions of the box?
Answer by ewatrrr(24785)

 
Hi,
2x^3+11x^2+17x+6 | Note IF x = -2, ⇒ -16 + 44 - 34 + 6 = 0
(x+2) one factor , could represent one of the three dimensions of the box